### v2

Abstract
We present Hapax Locks, a novel locking algorithm that is simple, enjoys constant-time arrival and unlock paths, provides FIFO admission order, and which is also space efficient and generates relatively little coherence traffic under contention in the common case. Hapax Locks offer performance (both latency and scalability) that is comparable with the best state of the art locks, while at the same time Hapax Locks impose fewer constraints and dependencies on the ambient runtime environment, making them particularly easy to integrate or retrofit into existing systems or under existing lock application programming interfaces Of particular note, no pointers shift or escape ownership between threads in our algorithm. 111source code:/https://github.com/davedice/Hapax-Locks.
keywords:
Synchronization; Locks; Mutual Exclusion; Mutex; Scalability; Cache-coherent Shared Memorycategory:
[…]
††margin: Context; Motivation; background; frame the problem; ††margin: endow; imbueLocks often have a crucial impact on the performance of parallel software, hence they remain the focus of intensive research with a steady stream of algorithms proposed over the last several decades. Picking an algorithm involves striking a balance between a number of factors, such as : latency in the uncontended case; throughput in the contended case; admission fairness; space requirements; coherence traffic induced by the lock; any dependencies the lock might have on other subsystems (such as dynamic memory allocation); and how well an algorithm might be retrofit under existing locking application interfaces. Our investigation focuses on algorithms and implementations thereof that gracefully tolerate large numbers of dynamically created and destroyed threads and locks, and that can be implemented easily under C++ locking interfaces and POSIX pthreads APIs. Applications expect to be able to acquire and hold multiple locks and to release in any order – scoped or balanced locking is not required and we have explicit lock and unlock operators. In addition, we prefer to avoid dynamic memory allocation, and the requirement for explicit constructor and destructor methods for threads and lock instances. In this paper, we present Hapax Locks, which possess all aforementioned properties.
2 Tidex Lock Algorithm
Our design starts with and is inspired by the Tidex[ppopp17-Ramalhete] mutual exclusion algorithm. Briefly, in Tidex, each lock instance has Arrive and Depart fields which contain thread identifier values. The lock is in unlocked state if and only if Arrive equals Depart. On arrival in the lock method, a thread atomically exchanges its own unique identity value into Arrive, which returns the predecessor value. The thread then waits for Depart to become equal to that predecessor value that it has in hand, at which point it has become the owner and can enter the critical section. In the corresponding unlock operator, the thread just stores its own thread identity value into Depart. That is, a thread in unlock announces that it has surrendered ownership by posting its own identity value into Depart – that outgoing thread never knows the identity of its successor, if any exists. Intuitively, when a thread is waiting and its predecessor departs and deposits its identity into Depart, that waiting thread becomes the next owner. Similar to CLH[craig-clh], the queue of waiting threads is implicit, with no manifest or explicit in-memory linked list. Waiting threads, if any, know only the identity of their immediate predecessor and wait for that value to appear in the Depart field. Collectively, however, a queue is formed. Identifier values are expected to be temporally unique and a given value can be associated with at most one thread at any given time.
A minor complication exists in that a thread could inadvertently try to acquire some lock , where had recently acquired and released , and ’s identity value was still installed in ’s Depart field. If were to simply atomically swap its identity value, , into ’s Arrive, then the lock would appear to be in unlocked state and we might suffer exclusion failure. The remedy, as described in [ppopp17-Ramalhete], is that each thread should maintain two identity values, a primary and an alternative. On arrival in lock, a thread first fetches Depart, and it if finds its own primary identity value there, left as a residual from prior operations, it shifts to its alternative ID for that lock-unlock episode.
| Time | Action | Owner | Arrive | Depart | Waiters |
| 1 | Initial unlocked state | - | 0 | 0 | |
| 2 | A arrives (enters) | A | A | 0 | |
| 3 | B arrives (waits) | A | B | 0 | B@A |
| 4 | C arrives (waits) | A | C | 0 | C@B B@A |
| 5 | A unlocks (pass to B) | B | C | A | C@B |
| 6 | B unlocks (pass to C) | C | C | B | |
| 7 | C unlocks (neutral) | - | C | C | |
| 8 | C locks (duplicate) | C | Ĉ | C | |
| 9 | C unlocks (duplicate) | - | Ĉ | Ĉ |
Table-1 depicts a scenario where threads acquire and release a Tidex lock. Initially, at time , the lock is in the unlocked state as the Arrive and Depart values are equal. For explication, we annotate the table with the identity of the current owner, but note that no such field is manifest in the lock data structure. At time , thread arrives and executes an atomic exchange operation to install into the Arrive field. The exchange returns . then fetches the Depart field and observes , and recognizes that it has acquired the lock without any need to wait. enters the critical section. At time , arrives to acquire the lock, and ’s exchange installs into Depart and returns . indicates that waits for to announce its eventual departure (unlock). At step thread arrives to acquire the lock. ’s exchange returns and waits for to announces its departure. In step , releases the lock by installing its own identity, , into the Depart field. As is waiting for to appear in Depart, has passed ownership to , and enters the critical section. In step , releases the lock, by storing into the Depart field, which is recognized by . is the new owner and enters the critical section. In step , releases the lock by storing into Depart. At this juncture both Arrive and Depart equal , so the lock reverts to unlocked state.
We continue the scenario to illustrate the requirement for alternative thread identities. At step 8, returns to re-acquire the lock. , however, also appears as a residual value in both Depart and Arrive. If were to simply exchange into Arrive, then the lock’s state would remain unchanged and we could suffer subsequent exclusion and safety failure. To avoid this scenario, arriving threads, before they execute the exchange into Arrive, must first fetch and inspect Depart, and, if their identity appears there, switch to an alternative identity, which, in this case, is .
Tidex is related to CLH[craig-clh] and HemLock[spaa21-Dice] in that contending threads wait on state related to the predecessor value that was returned from the atomic swap executed on arrival. The arrival “doorway” phase [cacm74-lamport] and unlock paths run in constant-time. While simple, the key downsides are the need on arrival to fetch and check the Depart field, and, most critical for scalability, the use of global spinning by waiters on the Depart field.
In Ticket locks[tocs91-MellorCrummey], in contrast, a thread waits for its own ticket value to appear in the lock’s Grant (“now serving”) field, whereas in Tidex a thread waits for its predecessor’s identity, obtained via the atomic exchange, to appear in a variable associated with the lock.
For reference, we provide an example Tidex Lock implementation in modern C++ in Listing-LABEL:Listing:Tidex. To assist the reader, we have annotated accesses to shared global locations in green. For brevity, this version expresses the critical section as a C++ lambda expression. Non-escaping lambdas are efficient and add no particular runtime overhead 222Example usage : Tidex L {}; int v = 5; L + [&]{ v += 2;} ;. The assert statements in the listing document invariants and do not constitute runtime checks against errant application usage of the lock. Critically, they are not required for correct operation and will be disabled in normal non-debug builds. To achieve the desired progress properties, we expect that C++ 64-bit std::atomic load, store, fetch_add and compare_and_swap run in constant-time, and, specifically, are not emulated via load-locked and store-conditional loops. We also assume the existence of a Pause() operator for polite busy waiting.
3 Hapax Locks
Hapax Locks are based on, but address the aforementioned performance shortcomings inherent in Tidex Locks. With Hapax Locks, threads no longer bother to maintain their own ID and alternative ID values. In lock, our thread notionally executes an atomic fetch_add(1) against a global shared 64-bit ID generator to conjure a new unique 64-bit value, and we use that as the thread ID for the acquire-release episode. The single generator instance is used for all threads and all locks in a process. The value from the generator is our hapax333In linguistics, hapax legomenon, – or just hapax (απαξ “once”) – refers to a word that appears just once in a body of text. In Hapax Locks, a particular hapax value is a nonce and appears at most once in the execution history of a process. value. Instead of a thread just having two Tidex-style identity values, it can now have multiple, allocated dynamically, one unique for each lock acquisition operation. This approach avoids any need to fetch and check Depart to ensure that our value does not happen to already appear in Arrive or Depart. A given hapax value, once allocated and made public by installation into Arrive, is never reused or recycled in the history of the process. With 64-bit values, we don’t, as a practical concern, need to worry about overflow and wrap-around of the hapax values generated. Unique hapax values are plentiful, and for all practical purposes, limitless.
To avoid the global hapax identity value generator becoming a coherence hot spot, in practice, we allocate blocks (consecutive runs) of unique hapax identifiers to a thread. Once it has allocated a block, our thread can then locally sub-allocate (just by incrementing the value) until the locally held block is exhausted, at which case it must reprovision from the global ID generator. The values generated from the local block are still globally unique. We argue this still avoids any practical real-world roll-over or wrap around concerns for the values. By reducing traffic on the ID generator, we avoid a scalability impediment. In particular, large blocks reduce the frequency of access to the global block allocator. In the attached code, we allocate via 64K blocks of hapax values, so the upper 48 bits of a hapax are associated with a specific thread and the lower 16-bit are used for local sub-allocation by that thread. A thread needs to reprovision from the global allocator every 64K lock acquisition operations. Developers can tune the block size as they see fit.
Instead of a using a single global allocator, we observe that more sophisticated and scalable allocators are possible. For instance, using an array of allocator “lanes”, offset appropriately, and selected randomly, or by geographic location in the system topology, will yield more scalable allocation, but such approaches were not found profitable as the allocation is amortized over the block size.
We note that threads might come into being, allocate a block, and then die, which, absent any other recovery mechanism, would cause all the remaining hapax values in that thread’s sub-block to be abandoned and discarded. In turn, such behavior might cause the global block allocator to overflow prematurely. As such, excessively large block values may not be appropriate. We believe a reasonable compromise to be 64K blocks with a 48+16 split.
Both Tidex and Ticket locks use global spinning, which can result in potentially excessive write invalidation rates 444Imagine a contended Ticket Lock with 10 waiting threads, and no arriving threads. When the owner surrenders the lock, the write will invalidate, at worst case, in 10 remote caches. Only one of those waiting threads that suffered invalidation will become the next owner and the remainder will continue to busy-wait, pulling the line back into their caches after the invalidation. When that new owner ultimately surrenders ownership, its write will invalidate in 9 remote caches, and so on. So if we have waiting threads, each with a private multiple-reader-exclusive-writer MESI-style cache, then the number of invalidations required to pass ownership to all the threads, assuming no new arrivals, is at least – potentially quadratic write invalidation rates. We refer to the number of caches that require invalidation from a given store instruction as the invalidation set or invalidation blast zone of that store. Broadly, larger invalidation sets may incur more latency and consume more interconnect bandwidth.. In Tidex, for instance, all waiting threads for a given lock busy wait on the Depart field. To avoid such global spinning, we take inspiration from TWA[EuroPar19-TWA, arxiv-TWA] – itself derived from ticket locks – and implement a shared waiting array, where all threads and all locks in a process use the same shared array. We opted to provision the array with 4096 slots. In theory, the number of slots should be a function of the number of logical processors in the system, but 4096 is reasonable even on very large modern systems. Threads perform long-term waiting via elements in the array. TWA hashed the ticket value and lock address to form an index into the waiting array, and the elements in those array slots were slot-specific sequence numbers. When TWA waiters observe the slot sequence number change, they would then recheck the actual ticket lock Grant “now serving” field to see if they were granted ownership. We refer to this as proxy waiting. Updates to the array elements constitute a conservative hint and notification to waiters that the actual location of interest may have been modified – the waiter is then expected to check the authoritative “ground truth” variable. In the TWA unlock operator, after updating Grant, TWA then atomically increments the corresponding slot sequence number to signal any waiters. Note, however, that because of collisions, waiters observing the sequence number change can not safely intuit that Grant had shifted to the waited-for value, so waiters are required to conservatively recheck Grant and, if necessary, resume the proxy waiting protocol. Absent hash collisions, we have private waiting, but as collisions may manifest, we say we have semi-private waiting.
For Hapax Locks, we could just apply any general hash function, taking the hapax value and, optionally, the lock address, to form an index into the waiting array. That would certainly suffice, but a better option is available. In the mapping function from hapax to waiting array index, we extract the high-order bits in the hapax, which are associated with the thread, and discard the rest (the offset into the thread’s block), and then apply a simple cache-aware lightweight hash to form the index. (We multiply by 17[WeylSequence] to map adjacent block zones onto different cache sectors or lines in the array, to reduce false sharing). So, for better performance, the hash — mapping function from waited-upon hapax value to waiting array index — is intentionally aware of the block allocation scheme. While not strictly required, we believe this is a useful optimization and strikes a balance between temporal cache locality (re-use by the same thread) in the array and the risk of collisions or near collisions (false sharing) between threads. This approach is illustrated in the ToSlot function in Listing-LABEL:Listing:Hapax-Locks.
In the unlock path, a thread writes its hapax — the one associated with that specific acquire-release episode – into Depart, and then also into the associated slot in the waiting array. This outgoing thread posts notification into the waiting array anticipating that its immediate successor, if any, will observe the change and recognize that it has been granted ownership. (If the successor observes the exact waited upon value appear in the slot, it can return expeditiously without needing to check Depart, otherwise if observes the value in the array change to any other value, it must conservatively recheck Depart). We require only simple C++ atomic loads and stores in the waiting array slots, unlike TWA, where updates needed to atomically increment the slot element. If desired, as an optional optimization, we can favor uncontended operation, and, in unlock, after storing into Depart, check if there are any newly arrived waiters, by fetching Arrive, and potentially skip the waiting slot update when there are known to be no waiting threads.
As noted above, if a waiter happens to see its waited-upon hapax value appear in the array, it does not need to recheck the Depart value to confirm handover, but, rather it can directly enter the critical section. We call this mode of transfer direct expedited handover. This technique is sound because the hapax values are globally unique and non-recurring. Crucially, this optimization could not be safely applied in TWA. A hapax value conceptually maps uniquely to a thread, a lock and a specific acquire episode by that thread on that lock. So in the normal case of no collisions in the array, we can effect a very fast direct handover where the recipient – the immediate successor – doesn’t need to check and confirm the Depart value. This, in turn, reduces coherence traffic, and makes handover recognition more efficient and reduces path latency. In the case of collisions in the array, a waiter might see the slot value change to a non waited-upon value. In that case it needs to recheck the actual Depart field value, to determine whether it is or is not the new owner. The waiter’s hapax value-of-interest could have flickered into visibility and then have been “stomped” or overwritten by unrelated concurrent actions, but not to the waited-upon value. So when a waiter observes the slot change to a value not of interest, it needs to recheck the definitive or authoritative “ground truth”, which is value in the lock’s Depart field. That scenario should be rare and only manifest via collisions in the array. Crucially, assuming hapax values never overflow and wrap around (which would compromise correctness) no ABA[ArtOf] pathologies are possibly given the non-recurring nature of hapax values – a slot will never flicker back to an old value, which, could potentially leave stranded waiters.
Under contention, the value of the current outgoing owner’s hapax variable, executing in unlock, is the same as the incoming immediate successor’s pred variable, as found in the successor’s lock phase. In Ticket locks, a thread knows that its immediate waiting successor must hold a ticket value of just one more than the ticket value allocated to the current owner. In Tidex and Hapax Locks, however, the departing thread has no particular knowledge about the successor, other than being able to determine if it exists.
In Listing-LABEL:Listing:Hapax-Locks we show an implementation of the Hapax Locks in modern C++.
4 Hapax Lock Variation with Visible Waiters
In Hapax Locks presented in Listing-LABEL:Listing:Hapax-Locks, above, when releasing a lock, the lock holder must store its hapax into both the Depart field, and then into the corresponding waiting array slot. (More precisely, the store into the slot must not become visible before the store into Depart). Each of these stores can incur a cache miss and prolong the lock handover transition duration, in turn limiting scalability[isca10-eyerman]. In this section, we describe an optimization that eliminates one of those stores in the common case, yielding an even more scalable variant. It achieves this by using visible waiters which in turn enables assured positive handover, which allows the thread executing in unlock to skip the store into Depart if it can confirm passage of ownership to its successor.
We say the baseline version of Hapax Locks shown in Listing-LABEL:Listing:Hapax-Locks uses invisible waiters as waiters only monitor the waiting array, but never write into the array, whereas in the version described in this section, with visible waiters, waiters explicitly announce their existence by updating slots in the waiting array.
In Listing-LABEL:Listing:Nonce-Visible-Waiters, accesses to shared globals in the contented fast path are annotated in purple and the key fast-path assured positive handover accesses are shown in blue. These reflect the dominant mode of execution under sustained contention. Critically, the atomic CAS (compare-and-swap) operation in unlock efficiently and expeditiously transfers ownership to its waiting successor running in lock.
By convention, we ensure a hapax value of 0 is never allocated. Normally slots in the array are 0 indicating they are unoccupied. Threads arriving to wait first try to CAS the slot associated with the waited-upon value (which is non-0) from 0 to that value. If that CAS fails, because the slot is occupied by another waiter, they then revert to degenerate Tidex-style global spinning. This should be rare and only occur on hash collisions in the array. Otherwise, that waiting thread owns and has reserved the slot for the duration of its waiting phase and then busy waits for the value to change to any other value. The corresponding unlock operator tries to clear the slot back to 0 with CAS. The slot value can then flicker to 0 and then to other values, because of concurrent activities, so the reader (waiter) is safe to interpret any change in the value as indicating that ownership has been conveyed. This protocol is safe by virtue of the hapax value non-recurring property. The slot can not ever revert to a previous value, so we do not miss wakeups to a waiter.
This variation effects contended lock ownership handover via 2 distinct modes : (a) The outgoing thread, executing in the unlock phase, uses an atomic compare-and-swap (CAS), on the associated slot in the array to attempt to overwrite its own hapax value with 0. The particular slot in the array is determined via a hash on the hapax value. Recall that the value of the current outgoing owner’s hapax variable, in unlock, is the same as the incoming successor’s pred variable in the successor’s lock phase. If the CAS is successful, the outgoing thread has accomplished expedited handover and foregoes updating the Depart field. The handoff is assured (positive) as, by virtue of the CAS, the departing thread knows with certainty that the waiter exists, and is waiting at the slot, and will be able to observe the change, providing a synchronous rendezvous between the departing thread and that successor. The waiting successor detects positive handover by observing that the hapax value it installed in the array, shifts, by virtue of the CAS by the predecessor, and can avoid checking the Depart field, and can directly enter the critical section. Such expedited positive handover – accelerated departure – is the preferred mode of succession, enjoying the shortest paths and least latency. Critically, under sustained contention, neither the outgoing thread, running in unlock, nor the incoming successor, waiting in lock, access any locations in the shared lock body; Hapax Locks shares this desirable attribute with MCS[tocs91-MellorCrummey] and CLH[craig-clh, ipss94-magnusson]. The outgoing thread, in unlock, does not need to store into Depart and the incoming successor, in lock, also does not need to fetch Depart to confirm handover, so we reduce coherence traffic on cache lines underlying the lock body. Unlock only needs to update the Depart field when there is no contention or if there were collisions in the waiting array. (b) The attempt at positive expedited handover fails – the CAS fails – so the outgoing thread reverts to the Tidex protocol and updates the Depart field, which is observed by the successor. Expedited positive handover can fail because of hash collisions on hapax values in the array, or if the unlock operation commences before the tardy successor has had a chance to register and make itself visible in the waiting array. The unlock operation can not determine with certainty if a successor exists – if the lock is contended or not – so it reverts to a more conservative protocol to release the lock.
With positive handover, unlock can also skip updating the Depart field during egress. if it can positively and definitively detect that it accomplished handover to a waiting successor. Furthermore, the successor can then also skip checking the Depart field. This optimization reduces write invalidations on the shared Depart variable. unlock only needs to update the Depart field when there is no contention or if there were collisions in the waiting array.
Visible waiters completely occupy a slot for the duration of the waiting phase, which means we might be more exposed to collisions if we have a large number of waiters. As such, a given slot can not be shared amongst multiple concurrent waiters. We note, however, that there is almost no particular downside to increasing the size of the array to statistically reduce the incidence of collisions. And we could, if desired, shift hapax zones if we detect collisions, in an agile and adaptive fashion, to reduce the odds of future collisions, providing a form of collision avoidance. Currently a thread falls back to simple Tidex global spinning if a waiter finds the slot occupied.
We also note that a thread could, as an optional optimization, operate briefly as a visible waiter, and if handover was not accomplished after a brief period, to avoid monopolizing an element, it could voluntarily vacate its slot and then shift back to the invisible waiting policy in the baseline algorithm.
The positive handover optimization enabled by visible waiters affords the most relative benefit under high flux situations, with high arrival and departure traffic rates on a given lock. If threads fail to acquire the lock during the brief initial visible spinning phase, then we know that throughput and progress over the contented lock is low, so the benefit of visible waiting and positive handover will not be as pronounced, so reverting to invisible waiting does not appreciably degrade performance in this particular operating regime. But by shifting to invisible waiting, our threads can reduce load on the visible waiting array, and allow other threads and locks to benefit from using positive handover via the waiting array.
Because of the possibility of collisions, the worst-case Remote Memory Reference (RMR) complexity [stoc08-attiya, podc02-anderson] for Hapax Locks is unbounded. If all hapax values were to inadvertently hash to just one singleton slot, then operation devolves to the underlying Tidex algorithm, with global spinning.
5 Related Work
††margin: semi-local; semi-private; quasi-private; commonly, frequently, often, usually mostly ::: fere- ; pene- ; quotide ; semi-; cotidie ; vulgo-; plerumque ; saepe-While mutual exclusion remains an active research topic [ppopp17-Ramalhete, craig-clh, ppopp91-Mellor-Crummey, EuroPar19-TWA, icdcn20-jayanti, Jayanti-Dissertation, EuroPar19-GCR, opodis17-dvir, EuroSys19-CNA, eurosys17-dice, oopsla99-agesen, topc15-dice, ScottB24, MCSH, isca10-eyerman, aksenov-CLH, aksenov-MCS, tocs19-Guerraoui] we focus on locks closely related to our design.
Simple test-and-set or polite test-and-test-and-set [ScottB24] locks are compact and exhibit excellent latency for uncontended operations, but fail to scale and may allow unfairness and even indefinite starvation. Ticket and Tidex locks are compact and FIFO, and also have excellent latency for uncontended operations, but they also fail to scale because of global spinning, although some variations attempt to overcome this obstacle, at the cost of increased space [EuroPar19-TWA, Ticket-AWN, dc03-Anderson, tpds90-Anderson, spaa11-dice]. For instance, Anderson’s array-based queueing lock [dc03-Anderson, tpds90-Anderson] is based on Ticket Locks but provides local spinning. It employs a waiting array for each lock instance, sized to ensure there is at least one array element for each potentially waiting thread, yielding a potentially large footprint. The maximum number of participating threads must be known in advance when initializing the lock. TWA[EuroPar19-TWA] is a variation on ticket locks that reduces the incidence of global spinning.
Queue-based locks such as MCS or CLH are FIFO and provide local spinning and are thus more scalable. MCS is used in the linux kernel in the low-level “qspinlock” construct [linux-locks, Long13, LWN2014]. Modern extensions of MCS edit the queue order to make the lock NUMA-Aware[EuroSys19-CNA]. MCS readily allows editing and re-ordering of the queue of waiting threads, [markatos, eurosys17-dice, EuroSys19-CNA] whereas editing the chain is more difficult under CLH, HemLock and Reciprocating Locks[arxiv-recipro, ppopp25-dice], where there are no explicit linked lists.
CLH is extremely simple, has excellent Remote Memory Reference (RMR) complexity[stoc08-attiya, podc02-anderson], and requires just a single atomic exchange operation in the Acquire operation and no atomic read-modify-write instructions in Release. Unfortunately the waiting elements migrate between threads, which may be inimical to performance in NUMA environments. CLH locks also require explicit constructors and destructors, which may be inconvenient or preclude their use in some environments 555Many lock implementations require just trivial constructors to set the lock fields to 0 or some constant, and trivial destructors, which do nothing. The Linux kernel spinlock_t/qspinlock_t interface provides a constructor, but does not even expose a destructor. Such decisions influence and may limit which lock algorithms can be readily implemented under a particular interface.. Similarly, C++ std::mutex is allowed to be trivially destructible, meaning storage occupied by trivially destructible objects may be reused without calling the destructor. Under both GCC g++ version 15 and Clang++ Version 20, is_trivially_destructible reports True for std::mutex, and as such, destructors do not run.
Our specific CLH implementation uses a variation on Scott’s [ScottB24] Figure 4.14, which converts the CLH lock to be context-free[ppopp16-wang], adhering to a simple programming interface that passes just the address of the lock, albeit at the cost of adding an extra field to the lock body to convey the address of the head waiting element to the corresponding Release. This field is protected by the lock itself.
The K42 [K42, ScottB24] variation of MCS can recover the queue element before returning from Acquire whereas classic MCS recovers the queue element in Release. That is, under K42, a queue element is needed only while waiting but not while the lock is held, and as such, queue elements can always be allocated on stack, if desired. While appealing, the paths are much more complex and touch more cache lines than the classic version, impacting performance. In addition, neither the K42 doorway nor the Release path operate in constant time.
HemLock combines aspects of both CLH and MCS to form a lock that has very simple waiting node memory lifecycle constraints, is context-free but still scales well in common usage scenarios. HemLock requires just a singleton waiting element (queue node) per thread, which can be placed in thread-local storage. HemLock does not provide constant remote memory reference (RMR) complexity [opodis17-dvir] in scenarios where a thread holds multiple contended locks. In this situation the single node suffers from multi-waiting[spaa21-Dice]. Similar to MCS, HemLock lacks a constant-time unlock operation, whereas the unlock operator for CLH and Ticket locks is constant-time. Unlike MCS, HemLock requires active synchronous back-and-forth communication in the unlock path between the outgoing thread and its successor to protect the lifecycle of the waiting element. We note, however, that HemLock remains constant-time in the Release operator to the point where ownership is conveyed to the successor. HemLock uses address-based transfer of ownership, writing the address of the lock instead of a boolean, differentiating it from MCS and CLH. Reciprocating Locks[arxiv-recipro, ppopp25-dice], like HemLock, requires just a singleton per-thread waiting element allocated in thread-local storage. Reciprocating Locks has constant-time arrival and unlock paths, but is not FIFO, instead providing 2-bounded-bypass.
We opted to exclude NUMA-aware locks such as Cohort Locks [topc15-dice, PPoPP12-dice] and Compact NUMA-Aware Locks (CNA) [EuroSys19-CNA] from consideration. We excluded Fissile Locks and GCR[EuroPar19-GCR] as they have lax time-based anti-starvation mechanism. Fissile, specifically, depends on the owner of the inner lock to make progress to monitor for starvation.
We also excluded MCSH[MCSH] which is a recent variation of MCS that uses on-stack allocation of queue nodes and hence supports a standard locking interface. Like MCS, the lock is not innately context-free, and additional information needs to be passed from Acquire to Release via extra fields in the lock body. In our experiments, the performance of MCSH is typically on par with MCS proper.
We further restrict our comparison to locks that use direct succession and hand off ownership directly from the owner to a specific successor and do not admit barging or pouncing where the lock is released even though waiting threads exist, which allow newly arriving threads to seize the lock.
| Lock Algorithm | |||||||||
| Property | MCS | CLH | HemLock | Ticket | TWA | Recipro | Tidex | Hapax | HapaxVW |
| Spinning | L+P | U+P | L+S | U+G | U+S | L+P | U+G | U+S | U+S |
| Constant-time Release | No | Yes | No | Yes | Yes | Yes | Yes | Yes | Yes |
| Context-free | No | No | Yes | Yes | Yes | No | No | No | No |
| FIFO | Yes | Yes | Yes | Yes | Yes | No | Yes | Yes | Yes |
| Path Complexity – Branches | 2+3 | 2+0 | 2+2 | 1+0 | 4+0 | 3+2 | 1+0 | 4+0 | 6+1 |
| On-Stack | No | No | No | N/A | N/A | Possible | N/A | N/A | N/A |
| Nodes Circulate | No | Yes | No | N/A | N/A | No | N/A | N/A | N/A |
| Explicit CTOR/DTOR Required | No | Yes | No | No | No | No | No | No | No |
| Thread CTOR/DTOR Required | Yes | Yes | No | No | No | No | No | No | No |
| Max Remote Misses per episode | 4 | 4 | 4 | 2 | 4 | 4 | |||
| Invalidations per episode | 6 | 5 | 5 | 10() | 8.5() | 4 | 10() | 5 | 4 |
| Dynamic Memory Allocation | Yes | Yes | No | No | No | No | No | No | No |
| Trivial TryLock | Yes | No | Yes | 64-bit | 64-bit | Yes | No | Yes | Yes |
| Space Complexity | |||||||||
In Table-2 we compare the attributes of various lock algorithms. Red-colored cells indicate potential undesirable properties. Note that all the locks provide a constant-time doorway phase. In the following we explain the meaning of each property.
Spinning : L indicates that waiters spin on a location that is usually collocated in the system topology with the waiting thread (typically on-stack or in thread-local storage) , which may confer performance benefits on some systems, such as those that use the Intel UPI[UPI] home-based coherence protocol, while U indicates that the location being waited upon is generally “unknown” and may be either local or remote to the waiting thread. P indicates private spinning, where at most one thread spins on a given location at a given time, G indicates global spinning, where all threads waiting on a given lock busy-wait on a single location, and S indicate semi-local[spaa21-Dice] spinning, where threads, absent hash collisions, enjoy private spinning.
Constant-time Unlock : indicates if the lock Release operation is bounded. As noted above, HemLock is constant-time up to the point where the lock is either released or transferred to a successor, but the Release operator, for reasons of memory safety, then waits for the successor to acknowledge transfer of ownership before the memory underlying the queue element can be potentially reused. Specifically, in HemLock an uncontended Release operation is constant-time and a contended Release is constant-time up to and including the point where ownership is conveyed to the successor.
FIFO : indicates the lock provides strict FIFO admission. The only non-FIFO lock is Reciprocating Locks, which still provides 2-bounded-bypass and starvation avoidance.
Context-free : indicates additional information does not need to be transferred from the Acquire operator to the corresponding Release operation.
Path Complexity - branches : We tally the number of conditional branches in the Acquire and Release methods, respectively, using the platform independent LLVM intermediate representation (IR) instructions, as emitted by clang++-20, which serves as a simple measure for path complexity. All Acquire operators have at least one branch, by virtue of the waiting loop. In some cases, we observed that the compiler converts if statements to conditional move operations, shifting control flow to data flow.
On-Stack : indicates the queue elements, if any, may be allocated on-stack. This also implies the nodes do not migrate and have a tenure constrained to the duration of the locking episode.
Nodes Circulate : queue elements migrate between threads. This often implies the need for an explicit queue element lifecycle management system and precludes convenient on-stack allocation of queue elements. Migration may also be unfriendly to performance in NUMA environments.
Explicit CTOR/DTOR Required : indicates the lock requires non-trivial constructors or destructors. CLH, when used in general purpose environments, for instance, requires destructors to run to release the wait elements referenced in the lock, to avoid memory leaks.
Thread CTOR/DTOR Required : indicates if per-thread constructors or destructors are needed to initialize or cleanup after dynamically created threads. We note that MCS implementations, could, if necessary, allocate queue nodes as needed, dynamically, in Acquire, and then release such queue nodes in Release. This avoids the need for thread-level destructors. (To avoid circular or recursive dependencies, we further require a specialized node allocator that is implemented via different types of locks). To avoid the allocator becoming a contention hotspot and bottleneck, common MCS implementations use per-thread caches of available elements, which need to be released when the thread terminates. Similarly, CLH could, in theory, allocate and free queue nodes on demand, although most implementations use a thread-local cache containing a singleton pointer to a free queue node, as seen in Scott’s [ScottB24] Figure 4.14. MCSH[MCSH] and K42[K42, ScottB24] MCS variations side step this concern by allocating elements on-stack, but at the cost of added complexity.
Max Remote Misses per episode : is the worst-case maximum number of misses to remote memory incurred, under simple sustained contention, by a matching Acquire-Release pair. Misses to remote memory – memory homed on a different NUMA node – may be more expensive than local misses on various platforms, such as modern Intel systems that use the UPI coherence fabric [UPI], where miss requests are first adjudicated by the home node of the cache line. Algorithms where nodes circulate are more vulnerable to accumulating such remote misses. For HemLock, we assume simple contention with no multi-waiting. We derived the “Max Remote Misses per episode” value via static inspection of the code, and determining which of the accesses that might cause coherence misses might also be to locations homed on remote nodes. We assume that back-to-back accesses of shared variables that reside on the same cache line, by a given thread, will result in just one miss. And for Hapax Locks we assume that the hapax value does not need to be replenished, which would incur an addition likely remote miss for the fetch-and-add atomic operation. We further assume no collisions (or near collisions, which could induce false sharing) in the waiting array.
Invalidations per episode : is the number of coherence misses, under sustained contention, experienced by an Acquire-Release episode in a given thread. We empirically approximate that number as follows. Using an ARMv8 system we modified the MutexBench microbenchmark (described below) to have a degenerate critical section that advanced only a thread local random number generator and to pass any context from Acquire to Release via thread-local storage, in order to reduce mutation of shared memory. Absent coherence misses, an Acquire-Release episode, including the critical section, manages to remain completely resident in the private L1 data cache. As such, any misses are coherence misses. This technique yields a useful empirical metric on how much coherence traffic various lock algorithms generate. We used the ARM l2d_cache_inval hardware performance counter, which tallies L2 cache invalidation events, as simple proxy for coherence traffic latency and bandwidth. We report the number of l2d_cache_inval events per episode. As can be seen, Reciprocating Locks and Hapax Locks with visible waiters are parsimonious and incur just 4 invalidations per episode, while CLH requires 5 666We observed similar ratios with performance counters that reported the number of “snoop” messages on the ARM coherent mesh “CMN” device, and on Intel by counting OFFCORE requests. We also observed in passing, while examining data from hardware performance counters, that CLH and MCS suffered from more stalls (both event counts and duration) from misses than did Reciprocating locks. CLH in particular executes a dependent load in the critical path in the arrival doorway on the address returned from the atomic exchange. The address to be loaded from is not known until after the exchange returns, denying the processor the ability to speculate or execute out-of-order.. Ticket locks and Tidex require (10 in our example) invalidations per episode, where is the number of participating threads, given the global spinning. The number of misses incurred by CLH, MCS, HemLock, Reciprocating Locks and Hapax Locks (assuming no hash collisions) is constant and not a function of the number of threads. These empirically-derived results align closely with a static analysis of the code and the expected number of coherence misses in the Acquire and Release paths. As the coherent interconnect fabric is a shared resource, excessive coherence traffic and invalidation rates can throttle scalability, and even impact the performance of other unrelated threads running on the system.
Dynamic Memory Allocation : indicates if the algorithm requires some type of dynamic memory allocation, typically for waiting elements. We note that MCS can avoid allocation and use on-stack allocation if all locking is lexically balanced and that CLH can also avoid allocation – by initially pairing elements 1:1 with locks and threads – if all threads and locks are statically allocated in advance of execution or where dynamically allocated threads and locks are effectively immortal. These advantageous circumstances are not normally found or available in general-purposes environments.
Trivial TryLock : indicates if the algorithm admits a trivial wait-free trylock operator. We note that Ticket-based locks can implement trylock if the Grant and Ticket fields are 64-bit bits, and effectively immune from roll-over overflow aliasing777Let’s say our ticket lock implementation has 32-bit Ticket and Grant fields. The naive but flawed way to implement trylock is to fetch the Grant field into a temporary , and then try to CAS Ticket from to . If execution is delayed between the load and the subsequent CAS, the values could have rolled around and we could suffer a false positive CAS success. This situation would be extremely rare, but not impossible. To avoid exclusion failure, if the CAS was successful, we might then try to adapt and remediate the algorithm to fetch and re-check Grant – after the successful CAS – and if Grant had changed, then we need to wait. This approach restricts us to 4G-1 possible concurrently waiting threads, but that’s arguably reasonable. But our trylock implementation is now flawed in the sense that it might need to wait, to recover. Another work-around for 32-bit fields is to collocate the Ticket and Grant fields within an atomic 64-bit superword and then to use a 64-bit CAS accordingly. This forces placement of the 2 fields, and can preclude attempts to limit false sharing by sequestering or isolating the constituent fields onto distinct cache lines. With 64-bit fields, however, given that rollover is not a practical concern, we can simply fetch Grant into and then try to CAS Ticket from to . The trylock attempt is successful if and only if the CAS was successful..
Space Complexity : reflects the space complexity of the lock algorithm where is the number of active threads, is the number of currently extant locks, is the number of locks currently held plus the number of threads currently waiting on locks, is the size of the waiting element. is the size of a lock instance. When a lock algorithm requires context to be passed from Acquire to the corresponding Release, we set to to indicate that our implementation allocated an extra word in the lock body to pass such information. Implementations that require context can avoid that space requirement in the lock body if they opt to pass context information by other means, such as via thread-local storage, in which case would be . HemLock is context-free, so the per-lock space usage is just , while Tidex, Ticket Locks, TWA and Hapax Locks require 2 words per lock. Hapax Locks and TWA require an additional 4096 words for the global waiting table that is shared over all thread and lock instances. For MCS and CLH we assume that the implementation stores the head of the chain – reflecting the current owner – in an additional field in the lock body, and thus the lock consists of head and tail fields, requiring 2 words in total.
6 Empirical Performance Results
We collected data on 2 different platforms. ARMv8 is an ARMv8 AARCH64 Ampere Altra Max NeoVerse-N1 non-NUMA system with 128 processors on a single socket, running Ubuntu 24.04. We compiled all code using the -mno-outline-atomics -march=armv8.2-a+lse flags in order to allow direct use of modern atomic exchange, CAS and fetch-and-add instructions instead of the legacy LL-SC (load-locked store-conditional) forms thereof. All lock busy-wait loops used the ARMv8 YIELD instruction. In the listings, the Pause operation reflects a single ARMv8 YIELD instruction. AMD is a 2-node system with AMD EPYC 9754 128-Core Processors (Ryzen) running Ubuntu 24.04. Each processor has 2 hyperthreads, yielding a total of 512 logical processors. The Pause operation is implemented as the PAUSE instruction.
Factory-provided system defaults were used in all cases. Unless otherwise noted, default free-range unbound threads were used, with no pinning of threads to processors or NUMA nodes. We used the default version of the GNU g++ compiler as provided by the Ubuntu distributions.
We compared MCS, CLH, TWA, Tidex, Ticket Locks, HemLock, Reciprocating Locks(Recipro), Hapax (HapaxIW), and Hapax augmented with the visible waiters optimization (HapaxVW).
6.1 std::atomic<>::exchange()
In Figure-1 we used a C++ benchmark, running on the ARMv8 system, that defines a simple structure type S that contains 5 32-bit int fields. The benchmark spawns concurrent threads, each of which loops, calling std::atomic<S>::exchange() to swap a local copy with a singleton global instance, followed by a non-critical section phase. We report the aggregate throughput rate, in completed exchange operations per second, at the end of a 10 second measurement interval, and we plot the median of 11 distinct runs in the figure. The number of threads is on the X-axis and aggregate throughput in the Y-axis. The C++ compiler and runtime implement std::atomic for such “large” objects by hashing the address of the instance into an array of pthread mutexes, and acquiring those as needed to implement the desired atomic action. Concurrent exchange operations on the same instance can thus result in contention on those underlying mutexes. We used the LD_PRELOAD facility to interpose on the pthread locking operations, replacing them with the algorithms of interest. Interposition allows us to change the underlying mutex without changing the application binary, by simply changing the LD_PRELOAD environment variable to reference various lock implementation libraries, which are loaded at process startup time. This allows for a fair comparision, but, like all pthreads implementations, precludes any compile-time optimizations that might otherwise be enabled at locking sites, between the application code and the locking algorithm. Any required lock context is collocated adjacent to the fields in the lock body.
In the non-critical phase, threads invoke a local xoroshiro128plus[Vigna-xoroshiro-PRNG] pseudo-random number generator (PRNG) to generate an integer uniformly distributed in [0,100) and then advance that same PRNG that many steps.
6.2 LevelDB
In Figure-2 we used the “readrandom” benchmark in LevelDB version 1.23
database888leveldb.org, again on the ARMv8 system, varying the number of
threads and reporting throughput
from the median of 5 runs of 50 seconds each.
Each thread loops, generating random keys and then tries to read the associated value from
the database.
We first populated a database999db_bench —-threads=1
––benchmarks=fillseq —-db=/tmp/db/
and then collected data101010db_bench —-threads=threads
—-benchmarks=readrandom
—-use_existing_db=1
—-db=/tmp/db/ —-duration=50.
We made a slight modification to the db_bench benchmarking
harness to allow runs with a fixed duration that reported aggregate throughput.
LevelDB uses coarse-grained locking, protecting the database with a single central mutex:
DBImpl::Mutex. Profiling indicates contention on that lock via leveldb::DBImpl::Get().
The results in Figure-2 largely echo the moderate contention results in Figure-3.
6.3 MutexBench benchmark
We also used a simple locking microbenchmark, MutexBench, to evaluate the performance of Hapax Locks. The MutexBench benchmark spawns concurrent threads. Each thread loops as follows: acquire a central lock ; execute a critical section; release ; execute a non-critical section.
For the moderate contention configuration, the critical section advances a shared global xoroshiro128plus[Vigna-xoroshiro-PRNG] pseudo-random number generator (PRNG) one step, and the non-critical section, as configured command-line arguments for this benchmark, advances a thread-local PRNG by 500 steps. In the maximum contention configuration, we set the non-critical section steps to be 0, yielding an empty non-critical section, subjecting the lock to extreme contention. At just one thread, this configuration also constitutes a useful benchmark for uncontended latency. At the end of a 10 second measurement interval the benchmark reports the total number of aggregate iterations of the outer loop (lock-unlock pairs) completed by all the threads.
We also report a trivial measure of long-term fairness, expressed as the number of iterations completed by the thread that made the least progress in the interval divided by the number of iterations completed by the thread that made the most progress. A value of is ideally fair and values toward less so.
After the measurement interval, the benchmark harness resets the global shared PRNG state and then advances, via sequential execution, the PRNG by the tally of steps taken over all the threads. As a cheap and racy test of exclusion and safety, we then ensure the shared PRNG re-arrives at the same state as observed at the end of the interval.
We repeated each data point 11 times, with independent runs, and reported the median throughput in Figure-3. For clarity and to convey the maximum amount of information to allow a comparison of the algorithms, the throughput on the -axis is offset from 0, and, as needed, the -axis is broken or log scale.
We note that HapaxVW Locks offer performance on-par with MCS, CLH, TWA, and HemLock, while performance falls either on-par or slightly under that of Reciprocating Locks. Tidex and Ticket locks fail to scale because of global spinning.
AMD-N1 reflects a configuration of the AMD system where constrained execution to just one NUMA node, with 256 processors (instead of the full complement of 512), in order to better examine NUMA performance effects111111numactl –cpunodebind 1 –membind 1.
7 Conclusion
††margin: claim, novel, novelty, contribution ††margin: Lesson, Informs, Instructs, Guides ††margin: Exemplar; explainer ††margin: MRAT = most recently arrived thread ††margin: strikes compromise; trade-off; balance; tension ††margin: derangement, disorder ††margin: Pick random starting position and direction ††margin: incur; suffer; induce; generate; require ††margin: KPIUnlike CLH, and like TWA, no pointers shift thread ownership in Hapax Locks (using the sense of ownership employed by the Rust and C++ programming languages [ownership], where responsibility for disposal of the referent transfers with the pointer), which is appealing as no memory lifecycle issues can exist, making it easier to reason about memory-safety. In HemLock, Reciprocating Locks, MCS and MCSH, threads spin privately and locally on memory that remains owned by the waiting thread, but the address of that memory escapes, and is exposed to other threads, which can read and write that memory. Under Hapax Locks, no thread-local (thread owned) addresses leak, escape, and are exposed and shared for access by other threads.
Hapax Locks, being value-based instead of address-based, avoid the lifecycle concerns present in CLH and MCS locks, as those lock algorithms must manage queue nodes. This property makes Hapax Locks particularly easy to integrate into existing systems, and in particular for low-level system software where the lock subsystem may not have access to facilities such as dynamic memory allocation. (Dynamic memory allocation facilities, such as malloc, are commonly implemented using locks, presenting a circular dependency. CLH and MCS typically allocate queue nodes via dynamic allocation, although the K42[K42] and MCSH variants are able to allocate nodes on-stack, and those nodes are required to be live only during the acquisition and waiting phases and not over the critical section or unlock operation. Hapax avoids such dependences and entanglements). Hapax Locks impose very few requirements or dependencies on the ambient runtime environment. Unlike CLH, Hapax Lock instances require only trivial constructors and destructors, and unlike both CLH and MCS, no allocation of queue nodes is required. As they never require allocation, or calls to any other services, the Hapax lock and unlock methods can be implemented as efficient leaf routines. And unlike Reciprocating Locks, Hapax Locks provide strict FIFO admission order.
Hapax Locks provide performance on par with existing state of the art algorithms, such a MCS and CLH. As it is value-based instead of address-based, as is the case for MCS and CLH, it avoids ensuant complications associated with allocating or deallocating memory for waiting elements, and concerns about the lifecycles thereof. The key benefit conferred by Hapax Locks is that they are easy to integrate or retrofit into existing systems, with minimal requirements, affording additional options for lock algorithm selection to systems developers.
Acknowledgements.
We thank Peter Buhr at the University of Waterloo for access to ARMv8 and AMD systems. We also thank the anonymous PPoPP reviewers for their helpful comments.References
Appendix A Key Properties of Hapax Locks
Strict FIFO-FCFS admission order.
Constant-time arrival and unlock paths.
Mitigates the multi-waiting issue found in HemLock.
Provides semi-private spinning, similar to that found in TWA. A lock provides private spinning if a most one thread busy waits on a given location at a given time and the write to that location invalidates the cache of just that waiting thread. Absent hash collisions in the waiting array, we have private spinning, but, given the possibility of collisions, we instead have semi-private spinning. As such, and because of the possibility of collisions, the worst-case Remote Memory Reference (RMR) complexity [stoc08-attiya, podc02-anderson] is unbounded. If all hapax values were to inadvertently hash to just one singleton slot, then operation devolves to the underlying Tidex algorithm, with global spinning.
Hapax locks do not provide guaranteed local spinning.
Space efficient : Hapax Locks require one word of thread-local storage per thread to track the current active private hapax block from which the thread is allocating; 2 words per lock instance for Arrive and Depart; one global word for the central hapax block allocator; and 4096 words – shared over all locks and threads – for the slots that form the waiting array. Any thread-local storage can immediately be reclaimed when the associated thread terminates. The waiting array is immortal.
Directly amenable to modern polite address-based waiting on elements in the waiting array. Examples of such techniques include WFE (ARM), MONITOR-MWAIT (Intel and AMD).
Nested and imbalanced locking are naturally tolerated and handled gracefully.
Appendix B Discussion
††margin: Unstructured bag of bullet items; middenTo preclude any potential concerns about roll-over in the hapax values, we expect implementations to use 64-bit hapax values, making roll-over a practical impossibility.
To ensure constant-time doorway and unlock phases, Hapax Locks requires that the hardware supports constant-time 64-bit atomic loads, stores, exchange, and fetch_add.
Like CLH, the algorithm requires only atomic loads, stores and exchange.
Unlike CLH, Hapax Locks can be trivially initialized with simple constants, avoiding the requirement of CLH dummy nodes. In fact Hapax Lock instances will normally be initialized to hold all values, which can confer an advantage for initialization, and also allows global lock instances to reside in the in the .BSS or .zero program segment instead of the .DATA segment. Relatedly, Hapax Locks do not require explicit destructors. In addition, and unlike common CLH and MCS implementations, which cache queue nodes in thread-local storage, no thread instance destructors are required. (We note that MCSH[MCSH] does not require such a thread-local cache as queue nodes can be transiently allocated on-stack).
As there are no CLH- or MCS-like queue node waiting elements, the algorithm is not vulnerable to queue node migration, which can prove detrimental on NUMA systems, for CLH.
TWA exhibits a tension on the waiting array size. Larger arrays reduce hash collisions, but, because of the undesirable “marching around the array” behavior, arising from the hash function and advancing ticket values, larger arrays could also increase cache footprint, cache pollution, and TLB misses. The access pattern in the waiting array induced by TWA can also trigger the processor’s automatic hardware stride-based prefetcher, which itself can induce false sharing and destructive interference. Relatedly, TWA provided no beneficial thread-level temporal re-use locality for accesses in the waiting array. In general, there is either no such trade-off, or a lessened trade-off with Hapax Locks. A bigger array is better, and reduces the odds of collisions or false sharing, although the relative benefits of a larger aray tends to fade as we increase size.
In Hapax Locks, however, locality in the array cuts both ways. For cache accesses, hash-induced locality acts to improve performance. But if we encounter collisions (or near proximal collisions that result in false sharing), then those collisions might persist for longer periods as a thread tends to re-access the same index until it reprovisions its block.
Assuming an “informed” hapax ToSlot hash function, a thread will exhibit strong temporal locality to a specific slot in the array, which, in terms of simple cache hit rates, is desirable.
Mixing the lock address into the hash function will tend to mitigate the effect of multi-waiting as found in HemLock. That is, Hapax Locks are less prone to multi-waiting than HemLock, although it can still occur because of collisions.
Hapax values are globally and temporally unique – one-time, one-shot, single-use, non-recurring, and disposable – and map to a thread and a lock instance and a specific acquire-release episode by that thread on that lock. Non-recurring values, which are ephemeral and evanescent, allow the algorithm to avoid ABA[ArtOf] pathologies. Once a hapax value is installed into any Arrive field, it is never reused in the lifetime of the process. The values are strictly unique over that tenure.
Once a hapax value is installed into any Arrive field, it is never reused in the lifetime of the process. The values are strictly unique over that tenure. For comparison, UUID[UUID, rfc9562] values, which are effectively 128-bit random numbers, are unique only with very high probability. A related concept is Snowflake IDs[SnowFlakeID], which incorporate a timestamp field and are also not strictly unique. We note that hapax values are effectively unordered, and only useful for equality comparison. (The current hapax allocation implementations happens to be ordered for a given intra-thread stream of values, in the sense that a larger value must have been allocated after a smaller value. Hapax Locks, however, do not depend on that particular implementation property.)
Hapax values are sub-allocated from thread-local blocks of consecutive hapax values, and thread-private blocks, when exhausted, are replenished by means of a global block allocator. Larger blocks mean reduced traffic on the global allocator.
In Ticket locks, for example, a thread waits for its own ticket value to appear in the lock’s Grant (“now serving”) field, whereas in Tidex and Hapax Locks, a thread waits for its predecessor’s identity, obtained via the atomic exchange, to appear in a variable associated with the lock.
The Arrive field contains the hapax identity of the most recently arrived thread while Depart holds the hapax identity of the most recently departed thread – the last thread to release the lock. The lock is in unlocked state iff Arrive equals Depart. The identity of the current owner is implicit.
KILL:REDUNDANTAs noted above, using hapaxes instead of fixed thread identities confers the following advantages. First, in the classic Tidex algorithms, arriving threads need to have at least 2 identity values, and must ensure that they don’t install a value into Arrive which happens to also reside in Depart as a residual value, otherwise the lock inadvertently appear to not be held, resulting in exclusion failure. To accommodate this requirement, Tidex, on arrival, fetches from the Depart field and switches to an alternative identifier, if the values matched. (Hapax Locks effectively provides an “almost infinite” stream of unique alternative identities). That fetch of Grant in Tidex adds to the path and coherence traffic generated by arriving threads. Second, threads waiting via the waiting array slots can recognize the specific identity of their predecessor, and effect handover without needing to consult the Depart field. In comparison, in TWA, which also uses proxy waiting with a waiting array, waiting threads, when they observe the sequence number in the waiting array change, must still ratify that they received ownership by rechecking the ticket lock Grant field against their own ticket number. This increases the handover latency and generates additional coherence traffic, compared to Hapax Locks. Finally, Hapax Locks can update the waiting array with just a simple store of the hapax value, whereas TWA must use an atomic operation to update the waiting array slot sequence numbers.
Compared to Tidex, Hapax Locks avoids the arrival redundant identity check, allows semi-private waiting in the array, whereas Tidex requires global spinning. Hapax Locks offers a direct handover optmistic path that is more efficient under contention.
As hapax values never repeat or reappear, we can use them instead of simple TWA-style sequence numbers – as change indicators – in the waiting array to indicate that the waiter may need to recheck the Depart field to confirm if ownership was conveyed to it.
For Hapax Locks with invisible waiters, in the case of hash collisions in the waiting array, a hapax value can be overwritten by a colliding access. In this case the thread waiting for one value might miss the direct handover optimization, but it will still observe the change and recheck the Depart value, recognizing handover in that fashion. Crucially, once a hapax value appears in the array, it will never reappear anywhere or at any time in the array. That is, hapax identity values are non-recurring and slot values thus never flicker back to a previous state.
The odds of collision in the array is equivalent to the so-called Birthday Paradox[BirthdayParadox]. ††margin: Equivalent to; Governed by Birthday Paradox
We observe collisions in the wait array can be either intra-lock or inter-thread.
Hapax Locks with invisible waiters can allow threads in different processes to safely synchronize on shared inter-process locks, as long as a shared singleton nonce allocator is used, and the waiting array is mapped into each address space. This is enabled by virtue of the fact that no addresses are shared between participating threads.
Hapax Locks are not context-free as we need to communicate an acquire-release episode’s hapax value – installed by the atomic exchange in the lock method – to the corresponding unlock method. We note that CLH, MCS, and Reciprocating locks all require context, while HemLock, Ticket Locks, and TWA are context-free.
In managed runtime environments such as the HotSpot Java Virtual Machine, with balanced locking under the “synchronized” construct, such context information can be kept in the stack frame. In C++, context can be kept in std::lock_guard or std::scoped_lock instance, which is typically allocated on stack. Rust’s Mutex<T> is similar. Such constructions are an example of the the Resource Acquisition is Initialization(RAII)[RAII] idiom. Context can also be conveyed to the unlock operation via a variant API, which allows passing of information, by using a lambda-based interface, through extra fields in the lock body, or via data structures held in thread-local storage that track held locks.
We would like to avoid collisions in the waiting array, but, for the purposes of performance, would also prefer to avoid near- or proximal-collisions in order to reduce spatial false sharing.
In sensitivity experiments with Hapax Locks with the number of ready threads set to more than the number of CPUs (intentionally forcing the system into preemption) and with the waiting array intentionally set to be small, we find significant rates of collision and sharing of slots in the waiting array. Many of the waiters are preempted, however, and not actively busy waiting. While collisions degrades performance because of the failure to provide private spinning, performance is actually dominated by the fact that we need kernel preemption to make forward progress, and preemption operates in geologic time. Assuming FIFO succession (or really, any form of succession that employs direct handoff), the choice of lock algorithm doesn’t play much of a role in performance in this regime of operation, when kernel preemption is in play. Performance abruptly falls off at the onset of preemption. There is not, for example, any appreciable performance difference between ticket locks (global spinning) and CLH (private spinning) at that point. The only locks that can withstand that kind of situation, without completely melting down, are those that use some type of bounded barging bypass or concurrency restriction, such as GCR, Fissile, etc. This tells us that we really only need to size the waiting array as a function of the number of the CPUs, and not the number of expected participating threads.
One downside to Hapax Locks is, like TWA, we can not guarantee that the locations being busy-waited upon in the array are homed on the same NUMA node as the waiting thread. In contrast, MCS and Reciprocating Locks provide locally homed spinning, where the spinning thread and the memory location being waited upon reside on the same NUMA node, which can yield a performance benefit on platforms such as Intel NUMA systems with the UPI[UPI, icpp23-katevenis] fabric which uses home-based coherence. In home-based coherence, all coherence probes on a given cache line are sent to the line’s home node, which adjudicates and arbitrates access. This approach can increase latency as more “hops” might be required, if the home needs to interrogate other nodes, but can reduce bandwith by reducing the need for broadcasts. On other common platforms, including Intel QPI, AMD, and ARM, performance is less sensitive to the location of the home node of a waited-upon location relative to the placement of the spinning thread. Instead, performance on those platforms, – coherence miss costs – is a function of the location of the requester, in the topology, the type of request (read or write), and the location of other caches that have the line, and the status of the line in those caches.
Similarly, and for the same underlying reason, TWA locks[EuroPar19-TWA, arxiv-TWA] perform well on all platforms except Intel NUMA UPI, where they lag behind other locks that use private locally homed spinning.
In Hapax Locks, the slot in the array where a thread waits for the predecessor’s value to be announced, signaling that the predecessor abdicated, is a function of the predecessor’s hapax value. As such, absent extra indirection, we do not have control where a thread waits – where the location in the array is homed. This is similar to CLH and HemLock, but unlike MCS and Reciprocating Locks, where waiting will be performed on locally homed locations.
Hapax Locks (with invisible waiters) provides dual handover encoding for succession on contended locks with waiting threads. The preferred mode is via expedited transfer via the waiting array while the fallback is via the Depart field.
In more detail, Hapax Locks effect contended lock ownership handover via 2 distinct modes in the unlock path. Recall that the outgoing thread, executing in unlock, updates the Depart field and then updates the associated waiting array slot with its own hapax values. If the waiting thread observes the waited-upon value appear in the slot in the array, it can exit the waiting phase immediately and enter the critical section, without any need to confirm the value of Depart. But if the slot in the array shifts to any other value, say , the waiter needs to recheck Depart. If Depart matches the waited-upon value, the thread has been granted ownership and can enter the critical section. Otherwise, the waiting thread will resuming waiting on the slot value and continue to wait while the value remains equal to . When the value next shifts, thread will again check for expedited handover, and otherwise recheck Depart, and, if necessary reenter the waiting loop using the recently observed value.
The waiting thread needs to triage 3 potential cases, while polling (busy-waiting) on the slot. If the value did not change since it was last fetched, the thread just continues to spin. If the value changed to waited-upon value, the waiting thread is now the owner, via direct expedited handover, and can enter the critical section. Othewise, the value in the slot shifted, so the waiting thread must conservatively recheck Depart before resuming the waiting protocol on the slot.
The particular associated slot in the array is determined via a hash on the hapax value, implemented in the ToSlot function.
Under contention, the value of the current outgoing owner’s hapax variable, executing in unlock, is the same as the incoming successor’s pred variable, as found in the successor’s lock phase.
Hapax Locks are thread-oblivious[topc15-dice] in that one thread can acquire a lock, but a different thread, in possession of the associated context, communicated from the acquiring thread, can subsequently release the lock. As such, they can be used in NUMA-aware Cohort Locks, as bounded binary semaphores, or other use cases where thread-oblivious locks are necessary. In addition, with the uptake of asynchronous futures, it is more common to find situations where one thread accquires a lock, and some other thread may subsequently release that lock 121212forexample:https://github.com/serenedb/serenedb/blob/main/libs/basics/unshackled_mutex.h.
In the code in the listings, we use 128-byte alignment and padding to reduce false sharing. The unit of cache coherence on Intel x86 is 64 bytes, but because of the hardware adjacent sector prefetch mechanism, we conservatively elect to use 128 bytes. We intentionally did not sequester or isolate Arrive and Depart as they tend to be accessed contemporaneously by arriving threads, exhibiting spatial locality. As such, the lock body a Hapax Lock is extremely compact.
Implementing try_lock is difficult if not impossible for Tidex. But under Hapax Locks, given non-recurring values, and the consequent elimination of ABA[ArtOf] pathologies, a thread in try_lock could inspect Arrive, and then Depart, and if the values were found equal, know with certainty that the lock is free, and the proceed to use an atomic compare-and-swap to try to install a new hapax value into Arrive, over the previous value. If the compare-and-swap was successful, the arriving thread has acquired the lock, otherwise, some other thread arrived and acquired the lock, and the try_lock operator can return the expected failure indication.
Relatedly, try_lock is difficult to implement in CLH and in ticket lock forms where the ticket field is only 32-bits and vulnerable to roll-over and inadvertent false-equal aliasing.
Consider implementing trylock in Tidex. We fetch and inspect the Arrive and Depart variables and find them equal, so the lock is apparently unlocked – apparently as we didn’t get an atomic snapshot of the two variables, as we used two loads. Ignoring the alternative ID issue, we then use CAS to try to swing Arrive to our thread’s identity. But the fields could have shifted underneath us, from other concurrent actions on the lock, yielding an ABA issue where our CAS was successful, but we then discover that Depart is changed, in which case we’d be forced to wait in trylock. We might also try to revert and back-out our update of Arrive, with another CAS, but there is no guarantee of success as other arrivals may have updated Arrive. The same scenario can manifest in ticket locks, although it takes 4G roll-over of 32-bit Ticket and Grant fields to occur. If we are willing to pack and collocate Arrive and Depart fields (or Ticket and Grant) fields together, and then use a double-wide CAS, you can safely implement trylock in that fashion. If we adapt ticket locks to use 64-bit fields, then, as roll-over and ABA are impossible, as a practical concern, we can safely implement trylock, where the operator first fetches the Grant field, observing value , and then tries to optimistically apply atomic compare-and-swap to swing the Ticket field from to .
Hapax Locks also gives us a viable trylock as we have 64-bits and non-recurring values. We further observe that trylock is safe even with the Hapax Locks visible waiters and positive handover. During positive handover episodes (sustained contention), we skip updating the Depart field (we resume updating Depart when contention subsequently abates) but we also know that Arrive and Depart must be different during such an episode, so the lock appears in locked state, and any trylock attempt will fail.
Like TWA – which is a variation of ticket locks augmented with a waiting array, to avoid global spinning – Hapax Locks are value-based and avoid passing shared addresses between threads. TWA and Hapax Locks also share similar RMR complexity by virtue of the shared waiting array. Unlike TWA, Hapax Locks provides expedited handover under contention.
In TWA and Ticket locks, a thread knows that its immediate waiting successor must hold ticket value of just one more than the ticket value allocated to the current owner. In Tidex and Hapax Locks, however, the departing thread has no particular knowledge about the successor, other than that it exists.
We could also employ futex address-based waiting, in the waiting array, but note that hapax values in the array must occupy 64 bits whereas the Linux kernel futex interface permits waiting only on 32-bit locations. Simply waiting on the low-order truncated 32-bits of a hapax is unsafe and could result in missed wakeups and progress failure. Instead, we could use the waiting chain construct in [TWASemaphore] where each slot in the array would be provisioned with an additional pointer to the head of a chain of waiting elements, implemented as concurrent pop-stack. Threads waiting on those chains would wait (using a futex) on a simple private flag variable in the elements they pushed onto the stack.
In the case of futex-based waiting, as we have semi-private waiting in the array, if a Hapax Lock is highly contended, then the set of futex addresses waited upon are distributed over the array, which, in turn, acts to diffuse kernel contention over the set of futex sleep channels, further improving scalability, as those sleep channels themselves are protected by kernel spin locks.
Hapax Locks are trivially amenable to using more sophisticated waiting techniques on the slots, such as WFE on ARMv8, MONITOR-MWAIT on Intel, or futex-based waiting. On ARMv8 systems we have observed that waiting via WFE makes a signficant improvement in contented throughput, compared to YIELD, for locks that uses private spinning, but that UMONITOR and MONITORX made no appreciable improvement for throughput on AMD and Intel processors.
We observe in passing that hapax values could be retrofit into the TWA locking algorithm. In unlock, instead of using an atomic fetch_add(1) in the waiting array slot, we could instead conjure a hapax value and simply store that value into the waiting array slot.
Hapax values could also be basis for counter-based random number generators[CBRN].
Hapax values can also be used in fashion similar to seqlocks[seqlock1, seqlock2]. Each record being protected by a hapax-seqlock has a payload area, a WriteFirst and WriteLast field, containing hapax values. To write the record, a thread creates a hapax value, stores that value into WriteFirst, then updates the payload, and finally stores the same hapax value into WriteLast. Note that writers, even when there are concurrent writers, never block or wait. Concurrent readers will first fetch WriteLast, fetch a tentative copy of the payload, and then fetch WriteFirst. Critically, readers fetch the hapax fields in reverse order relative to the write order. If the two fetched values match, then the reader has confirmed that the payload fetched is a consistent atomic snapshot, and is safe to use. Otherwise the snapshot is inconsisent, and the reader takes action to respond accordingly. The snapshot can be inconsistent because at least one concurrent write is in progress, or when previous multiple concurrent writers overlapped and interleaved accesses, leaving the record in an inconsistent state. This technique is useful for concurrent caches, for instance, where inconsisent values can be reconstituted and re-installed, recovering from the inconsistent state. We expect the payload to be trivially copyable, overwritable without risk of leaks, and to not require destructors, as the value can be “stomped” by concurrent writers. We observe that using this technique in modern C++ requires particular care[SeqLockMemcpy].
Many platforms exhibit topological favoritism or, specifically, NUMA favoritism, where atomic operations that are either near (in the topology) the home node of the cache line or are near the cache that currently holds the line in modified state are more prone than far requesters to be successful and serviced in the near time. Perversely, test-and-set locks are NUMA friendly on such platforms in that they reduce lock migration[topc15-dice, PPoPP12-dice, arxiv-CNA, EuroSys19-CNA] by virtue of such architectural unfairness, offering improved performance compared to pure FIFO locks. Backoff mechanisms commonly used in test-and-set locks also exacerbate unfairness, as a waiting /thread that has backed off is less like to acquire the lock in unit time than a more recently arrived lock. Backoff also represents “dead time” and is not work-conserving, as it is possible, under varying load, that the last waiting thread is stalled in a long backoff period when it could otherwise acquire the lock and drive progress.
We believe it is safe to aggressively fetch Depart early in the lock operation, before performing the atomic exchange on Arrive. That is, we can lift or hoist the fetch of Depart, which normally apears after the exchange, to before and “above” the exchange. Such an optimization, with early fetch of the Grant field, is also safe in Ticket Locks if we use 64-bit Ticket and Grant fields.
We call out a confounding benchmark phenomenon we call the re-arrival effect. We say a lock algorithm is unlock-terse if, under sustained contention, it avoids updating lock body fields in unlock, and unlock-prolix if it usually updates the lock body in unlock. Unlock-prolix lock algorithms that always write into the shared lock body, in unlock, such was Ticket Locks, will usually suffer more coherence traffic than do unlock-terse algorithms – such as MCS and CLH – that avoid, under sustained contention, updating the global lock body state in unlock.
Say we have a maximum contention benchmark with an empty non-critical sections, and sustained contention. Threads release the lock in unlock and then immediately re-arrive in lock to re-acquire the same lock. But, with an empty critical section, and assuming all the lock body fields are collocated on the same cache line, the updates by an unlock-prolix algorithm in unlock will drive the line(s) underlying the lock body into M-state, and the subsequent lock operation will re-access those same lines while they still reside in the local cache in M-state, as no other lock operations have had an opportunity to arrive or depart and invalidate the line. This effect somewhat distorts performance and confers a relative performance benefit to prolix lock algorithms in the case of maxmimum sustained contention with no intervening non-critical section. The degree of impact is specific to the architecture and the lock algoirthm – whether prolix or not.
The lock operator, for an unlock-prolix algorithm, can avoid coherence misses that it might otherwise incur if threads arrived between the lock and unlock.
We note that the existence of lock context passed through the lock body can trigger the re-arrival effect, even if the algorithm is normally unlock-terse.
Using randomized non-critical section durations acts to mitigate this effect, and is also arguably more faithful to the real-world application behavior we are trying to model in benchmarks.
Similar to CLH and Reciprocating locks, no explicit queue of waiting threads is made manifest. Waiting threads know only the identity of their immediate successors. As such, these locks can not trivialy be adapted, when implemented within a pthread mutex-condvar environment, to enable the classic wait morphing optimization, where signaling a condition variable within a critical section simply moves threads from the condition variable’s wait set to the mutex’s queue of waiting threads.
Another benchmarking pitfall is as follows. Say we have a FIFO lock algorithm, and, for the purposes of this explanation, an empty critical section. It it then common to find a long-standing repeating cyclic admission schedule. Often, this schedule is set via happenstance at startup, but then persists for long periods. A given schedule can be either favorable or unfavorable, for throughput, in terms of the system topology. Using NUMA as an example, a given schedule might reflect either many or few lock migrations, which in turn will reflect in the throughput achieved by that schedule. But other topology structures, such as AMD’s CCX clusters, can yield similar favoratism. This, in turn, can result in higher run-to-run variance, if we assume the initial schedule arrises in a mostly random fashion. This effect can be mitigated in part by injecting randomized non-critical section delays, in order to shuffle or shift the schedule and avoid long-standing persistent schedules.
TWA vs Hapax. Naively, we might expect the baseline Hapax Locks (with invisible waiters – the baseline algorithm) to yield about the same performance as TWA. But Hapax often underperform TWA for the following reasons:
For the hapax forms, we need to pass the hapax value we installed, as context from a lock operation to the corresponding unlock() operation. In the LD_PRELOAD implementations, we keep any lock context as an extra field in the lock body, but unfortunately accessing the context field generates coherence traffic. But for TWA (like ticket locks), we don’t need to pass context, but can instead just increment (non-atomically) the Grant field in the lock body.
TWA enables an optimization where we let the thread that’s the immediate successor shift from waiting in the waiting array to direct spinning on the Grant field in the lock body, hopefully yielding a more expeditious handover. Whether this specific optimization is profitable or not is very platform-specific and related to coherence costs. The ticket value a thread has in-hand gives a sense of proximity to the front of the queue. There’s no such optimization available for the Hapax family. The hapax value doesn’t really convey any sense of proximity to the “front” position in the logical queue of waiting threads, whereas a ticket value does.
TWA, however, exhibits the “marching around the waiting array” effect, which is unfriendly to caches.
Appendix C Optional Optimizations
††margin: Optional optimistic optimizations opportunities ; embellishments; variationsWe can intuit thread-vs-thread collisions when a waiting thread observes the value in the waiting array shift, but not to the waited-upon value. In this case, to reduce the odds of further collisions at that same waiting array index, the thread could optionally abandon its current block of hapax values, which are known to collide, and, the next time it allocates a hapax, allocate an entirely new block. While this technique for collision reduction is viable, it consumes hapax values at a higher rate and might eventually result in premature hapax roll-over. As an option, a thread might hold 2 active blocks and switch between them based on collisions.
There are some other approaches we could use to reduce waiting array collisions. One thought would be to just use an arbitrary “proper” hash that’s not hapax-aware. Intra-thread temporal access locality in the array is much reduced. Another idea would be to hash on the lock address, the thread ID part of the hapax, and the sub-sequence part of the hapax, but throwing away, say, the lower 5 bits of the local sequence number. So a thread would go to the same slot in the array 32 times in a row, which should still confer some locality benefits, but any specific thread-vs-thread collision scenarios would persist for only 32 episodes. This constitutes a reasonable compromise between collision penalties and cache locality benefit.
To further allay any residual concerns about hapax overflow we might shift to a hapax formulation that has a 64-bit thread ID field and a 64-bit sub-block value. At that point all concerns about overflow are eliminated. Modern Intel x86 and AMD processors now guarantee that properly aligned 128 byte loads and stores are completely atomic via the movdqa instruction. Using 128-bit hapax values would increase the lock instance size. Also, the Intel instruction set only has lock:cmpxchg16b for 128 byte atomics. There is no 128-byte atomic exchange. So we’d need to emulate the exchange with a CAS loop, and we lose our constant-time arrival doorway step and are potentially more vulnerable to architectural induced unfairness. Modern ARM also has 128-byte loads and stores and LL-SC. Encoding tuples in this fashion provides a safe solution that is immune to roll-over but which is not viable commonly available current processors.
The following variant of the Hapax algorithm works well in a managed runtime environment with garbage collection, like that provided by the HotSpot Java Virtual Machine. Instead of conjuring up our hapax values, we just use a reference to a freshly instantiated object as the hapax identity : hapax = new Object(). These objects will typically be allocated out of the thread-local allocation buffers (TLABs) via a bump pointer, so in the normal case such allocation is extremely fast and requires no synchronization. The astute reader may notice the analogy between hapax blocks and TLABs. Garbage collection also ensures the hapax values are temporally unique, even in a 32-bit JVM. Of course this technique offloads some of the problem off onto the garbage collector and object allocator. Allocating hapax values consumes space, which must eventually be recovered. The arrival path is no longer constant-time. In addition to possibly needing locks, it could in theory even induce a full garbage collection. Overflow is impossible, though.
Depending on the hapax allocation scheme and hash function, which maps hapax values to waiting array indices, if a given block, or portion of a block, maps to just one slot, then we can precompute and “memoize” the slot index and store the value in thread-local storage and compute the hash just once per block.
That is, depending on the block size, and specifics of the hash function, in ToSlot, which maps hapax values to indices (slots) in the waiting array, a thread may be able to precompute and cache in thread-local storage – or “memoize” – the address of the waiting array slot it will use for a sequence of hapax values, avoiding some computation.
Given that uncontended locking is extremely common, to reduce the consumption and burn date of hapax values, we might institute optimizations to recover values from uncontended locks, where those values never leaked or were exposed to other threads. A thread could maintain a simple cache of hapax values it recovered from re-locking the same lock instance. While workable, we do not currently believe the bookkeeping overheads associated with this specific optimization make the technique profitable.
If we expect contention to be rare, then we can apply an optional optimization in the unlock path. After storing in the Depart field, a thread will then fetch from Arrive. If the values are the same then we have reverted to unlocked state and can safely elide the store into the waiting array. This optimization reflects something of a bet or trade-off. If we try to apply the optimization and fetch from Arrive but find contention (the fetched value differs from the value we just stored into Depart) then we still need to store into the waiting array, but have generated additional futile coherence traffic by fetching from Arrive, to no avail.
A reasonable heuristic is a follows. If, when we acquired the lock, our thread had to wait, then we can assume prior contention on a lock accurately predicts future near-term contention. In that case, we bet that contention still exists, and update the waiting array unconditionally. But if we acquired without waiting then, in the corresponding unlock operation, we try to apply the optimization and condition the store into the array based on the value fetched from the Arrive field.
Listing-LABEL:Listing:Nonce-Allocator describes a variety of optimizations that can be applied to the hapax block allocation procedures. Some of the techniques therein might be useful in environments where thread-local storage is not available or inefficient, and hapax values would need to be allocated one-at-a-time.
The visible waiters Hapax Locks form can optionally optimize uncontended unlock operations as follows. If, in unlock, we find that Arrive still equals the hapax value installed by our thread, in lock, then there are no threads waiting on the lock, and we have a simple uncontended unlock operation, so we can safely use CAS to try to attempt to swing Arrive back to the value found in Depart, If the CAS is successful we have rolled back the state to unlocked and can then return without any need to update the waiting array slot. That is, instead of updating Depart, we roll-back and revert Arrive. Furthermore, if desired, as that hapax value was thus never observed by other threads, we can reuse it for subsequent operations. ††margin: Revert; roll-back
Appendix D Optimized Hapax Allocator Variations
We could, as an optional optimization, provide a small array of global ID generators, initialized with properly skewed values to ensure uniqueness, and when we need to reprovision, pick one “lane” at random and allocate from that element. This diffuses coherence activity over the array and avoids the situation where the allocator might itself constitute a hot-spot. In turn, this would make smaller block sizes more practical, reducing the cost a thread needs to reprovision. There doesn’t seem to be any practical benefit for this optimization, however, if the block size is sufficiently large to amortize the cost of allocating new blocks from a singleton allocator, as it will be accessed infrequently. Listing-LABEL:Listing:Nonce-Allocator shows an implementation of such a hapax allocator.
Instead of randomly picked slots, another viable method is to arrange the array of generators to be indexed with the CPUID or NUMA node ID of the allocating thread. This greatly reduces coherence traffic, assuming each generator is properly sequestered on its own cache line or sector. Interestingly, such hapax values also convey geographic position in the system topology – via the CPUID or NUMA node ID – of the associated thread,
Appendix E Hapax Lock Variation with Visible Waiters
Listing-LABEL:Listing:Nonce-Visible-Waiters-FULL reflects a variation that uses visible waiters, which in turn enables assured positive handover. This particular expanded listing contains augmented comments and optimizations.
Appendix F Additional discussion regarding the Visible Waiters variant
††margin: Dominant hot pathUndesirable aspects of this variant, compared to baseline Hapax Locks, include the following :
The visible waiters variant suffers from more complex paths, as there are races we need to recover from when threads start waiting at about the same time as their signaling unlock occurs.
††margin: Coherence out-of-order “transaction” width; counted in accesss; arbitration window; group of accesses covered by just once coherence transaction; coalesced coherence operations; atomicity granularity; Miss shadow; grace period; We posit and conjecture the existence of a small post coherence miss retention interval where the requestor can hold the line against remote arbitration for a short period. The existence is not guaranteed and is likely an implementation artifact, but targetted ad-hoc experiments suggest it exists
There are additional accesses to recheck shared globals, but those accesses tend to be to locations just accessed a few instructions ago and which are likely still resident in our thread’s local cache. While our approach exhibits good spatial and temporal locality, the benefits of that residency, in terms of coherence traffic, are timing-, platform- and application-dependent.
The approach is relatively atomic-heavy and employs a number of CAS instructions. We note that counting atomics (CAS), once popular in concurrency literature – and, at the time, relevant, given that legacy atomics were implemented via locking the bus – may be somewhat anachronistic, given that modern processors (AMD Ryzen, apple M series) implement CAS locally, in cache, have much reduced the local latency and pipeline disruption induced by a CAS. Arguably, CAS isn’t much different than a fenced store, and in particular the main cost of a CAS is, like a store, any necessary coherence-related write invalidation.
Waiters completely occupy a slot for the duration of the waiting phase, which means we might be more exposed to collisions if we have a large number of waiters. As such, slots can’t be shared amongst multiple concurrent waiters. We note, however, that there is almost no particular downside to increasing the size of the array to statistically reduce the incidence of collisions. And we could, if desired, shift hapax zones if we detect collisions, in an agile and adaptive fashion, to reduce the odds of future collisions, providing a form of collision avoidance. Currently the code falls back to simple Tidex global spinning if a waiter finds the slot occupied, but we could be more refined and add an extra field to the slot, and just fall back to classic Hapax Locks (without visible waiters) in the event of collisions between waiters. So a slot would then have at most one “privileged” visible waiter but could support multiple invisible waiters in the usual Hapax Lock fashion.
The uncontended unlock path is more complex, but in practice it doesn’t seem to make any performance difference as the path is just re-accessing locations that it already just touched.
Desirable aspects :
We provide expedited handover under contention. The critical paths in unlock and on the waiting side are short with few global accesses.
Unlock can also skip updating the Depart (Egress) global if it can positively detect that it accomplished handover to a waiting successor. The positive handover property lets us safely skip the update of Depart in the unlock operation. This optimizations reduces write invalidations on the shared Depart variable. unlock only needs to update the Depart field when there is no contention or if there were collisions in the waiting array.
In experiments with Hapax Locks with the number of threads set to more than the number of CPUs (intentionally forcing the system into preemption) and with the waiting array intentionally set to be small, we find lots of preempted waiters occupying the waiting array slots which in turn forces other waiters to revert to the cruder global spinning waiting mode. As noted above, however, performance is dominated by the affects of preemption and lock design, and array collisions are not of particular importance. And, reiterating, the key tunable parameter for sizing the waiting array is the number of physical CPUs.
The only exception to that claim might be if the system is using a non-standard non-1:1 threading model, where, say, we had 1000 logical runnable virtual threads, but they were multiplexed over a smaller number of real kernel threads. In that case we could probably end up saturating the waiting array and actually see some performance fall off because of array collisions and occupancy. But at that point we would also been required to modify the locking subsystems to use virtual thread-aware parking mechanisms, instead of spinning or traditional parking, and the new parking mechanism would have be codesigned with the user-mode thread model and be able to voluntarily switch, in user mode, off the stalled virtual thread (waiting for the Hapax Lock) to some other runnable virtual thread. As such, Hapax Locks might not be an appropriate choice for environments that use non-1:1 user mode threading models.
We observe, however, that Hapax Locks variant in Listing-LABEL:Listing:Nonce-Visible-Waiters could be made threading-model aware, if, when waiting, immediately before they dismount from and abdicate a kernel thread (switching that kernel thread to carry another runnable 131313A runnable thread is ready and eligible for dispatch, but is not currently running on kernel thread user-mode virtual) they also surrrender their slot in the waiting array in order to be altruistic and “polite” and decrease the odds of collisions in the array for other running threads. When the virtual thread is eventually dispatched and re-mounted on a kernel thread, it would then recheck Depart, and, if necessary, try to reclaim the slot and reinstall itself in the array.
In order to reduce occupancy pressure in the global waiting array, we can optionally institute a 2-stage waiting strategy where an arriving thread will first busy-wait as a visible waiter, but after some period, the thread will then voluntarily vacate its slot in the global array, to avoid monopolizing the element, and switch to a longer-term waiting mode that uses invisible waiters, such as found in baseline Hapax locks.
That is, threads busy-wait briefly as visible waiters via the waiting array, but, if they have not been granted the lock within a short period, they remove their visible waiting marker from the array, surrender the slot, (retracting the visible waiting state indication) and revert to longer-term invisible waiting in a second global shared waiting array, which runs parallel to the primary visible waiting array. There, threads employ a second spinning phase, as invisible waiters, followed, if desired, by futex-based waiting on the associated array element.
The positive handover optimization enabled by visible waiters affords the most relative benefit under high flux situations, with high arrival and departure traffic rates on given lock. If threads fail to acquire the lock during the brief initial visible spinning phase, then we know that throughput and progress over the contented lock is low, so the benefit of visible waiting and positive handover will not be as pronounced, so reverting to invisible waiting does not appreciably degrade performance in this particular operating regime. But by shifting to invisible waiting, our threads can reduce load on the visible waiting array, and allow other threads and locks to benefit from using positive handover via the primary array.
Each slot in the array could be provisioned with one field for exclusive visible waiting – where the visible waiter registers itself – and another field configured in the fashion of baseline Hapax locks, which allows multiple concurrent invisible waiters (arising from collisions) to monitor the slot. This allows the slot to be used in both modes. Alternatively, we could use 2 distinct arrays, running in parallel, one for visible waiting and another for invisible waiting. In addition, if a thread arrives and tries to register but finds the visible waiter element occupied, it can fall back immediately to baseline invisible waiting, instead of reverting to degenerate global spinning on Depart.
As noted above, the visible waiting technique is preferred in the case of high flux locking, with high arrival and traffic rates, as the fast positive handover property provided by visible waiting confers benefits for throughput over the contented lock, allowing ownership to be conveyed more efficiently and with less latency. Using invisible waiters does not permit the fast positive, handover, but if a thread has waited through its short-term grace period and switched from visible to invisible waiting, it is likely that any reduction in performance from using invisible waiting is relatively negligible as the lock is not enduring high throughput rates.
Appendix G Expanded Hapax Lock listing – with Invisible Waiters
Listing-LABEL:Listing:Nonce-InvisibleWaiters-FULL is an expanded listing augmented with additional comments and optimizations.